OpenStudy (anonymous):

WILL MEDAL AND FAN!! NEEDS THE ANSWER Write 12/20 as a percentage.

OpenStudy (anonymous):

Just divde 12/20 and times it by 100

OpenStudy (anonymous):

ugh

OpenStudy (anonymous):

ik how to do it, but i'm in a hurry

OpenStudy (anonymous):

Well then use a calculator.... Very simple.

OpenStudy (anonymous):

here is a great calculator to us https://www.desmos.com/calculator

OpenStudy (anonymous):

i have a phone

OpenStudy (anonymous):

Ok, use your phone. That works too.

OpenStudy (anonymous):

1 is the answer to the division

OpenStudy (anonymous):

you can also get the answer by multiplying the top by 5

OpenStudy (anonymous):

look i need the answer. anyone please help

OpenStudy (anonymous):

12/ 20 x 100

OpenStudy (anonymous):

\[\frac{12}{20}=\frac{5\times 12}{100}\]

OpenStudy (anonymous):

You said you know how to do it so all you need is your phone?? Why do you need us to give you the answer?

OpenStudy (anonymous):

\( \huge \frac{12}{20} \times 100 \)

OpenStudy (anonymous):

why does it matter to you? i said I NEED THE ANSWER

OpenStudy (anonymous):

Have fun.